OpenStudy (anonymous):

Solve: 6x2 + 12x = 0 (Points : 4) {–6, –2} {0, 2} {0, –2}

OpenStudy (tanya123):

6x^2 + 12x = 0 Factor: 6x(x+2)=0 x=0 OR x+2=0 x=0 OR x=-2